how many moles of chlorine gas would occupy a volume of 35.5L at a pressure of 100.0 kpa and a temperature of 100. c? after determining the number of moles calculate the number of grams of chlorine (cl2) contained in this container


1
Expert's answer
2021-11-18T10:01:03-0500

V= 35.5 L

p = 100.0 kpa = 0.9869 atm

T = 100 +273 = 373 K

Ideal gas law

pV = nRT

R = 0.08206 L×atm/mol×K

"n= \frac{pV}{RT} \\

n = \frac{0.9869 \times 35.5}{0.08206 \times 373} = 1.144 \;mol"

M(Cl2) = 70.9 g/mol

"m = 1.144 \times 70.9 = 81.11 \; g"

Answer: 1.144 mol and 81.11 g